what is the equation for the line that passes through the points (3,5)and(4,1)?
Answer by checkley75(3666)

FIRST WE FIND THE SLOPE
SLOPE(m)=(Y2-Y1)/(X2-X1)
m=(1-5)/(4-3)
m=-4/1=-4 THIS IS THE SLOPE.
NOW WE SUBSTITUTE ONE SET OF X,Y POINTS FOR THE X & Y IN THE LINE EQUATION [Y=mX+b] & SOLVE FOR THE Y INTERCEPT (b)
5=3*-4+b
5=-12+b
b=5+12
b=17 ANSWER FOR THE Y INTERCEPT.
Y=-4X+17